# ECAFormer: Low-light Image Enhancement using Cross Attention

### v2

@@ Introduction
Experimental Configurations
Datasets
@@
We conducted comparisons across multiple public datasets: LOL-v1 (Wei et al. 2018), LOL-v2 (Yang et al. 2021b), SID (Chen et al. 2019), SMID (Chen et al. 2018), and SDSD (Wang et al. 2021). Additionally, we have collected the Traffic-297 dataset, which includes various traffic scenes. It contains an amount of real nighttime traffic scenes with complex lighting environments, which pose a significant challenge to low-light enhancement models.
Explementation Details
@@
We implemented the ECAFormer model using PyTorch. The model was trained on an NVIDIA RTX 4090 with 24 GB of VRAM using the Adam optimizer ($\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999), across a total of 250,000 iterations. The initial learning rate was set at 2 $\times$ 10-4 and gradually reduced to 1 $\times$ 10-6 through a cosine annealing schedule during the training process. Training samples were generated by randomly cropping 128 $\times$ 128 patches from pairs of low-/normal-light images. The batch size was set to 8. The training data was augmented with random rotations and flips to enhance variability and robustness. The training objective was to minimize both $\mathcal{L}_{p}$ Perceptual Loss and $\mathcal{L}_{c}$ Charbonnier Loss between the enhanced images and their corresponding ground truths, ensuring high fidelity in the restoration of image details and color accuracy. To evaluate the performance, we employ the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) (Wang et al. 2004) as the primary metrics for evaluation.
Comparison Methods
@@
To validate the effectiveness of our derived ECAFormer, we conducted extensive comparisons with several state-of-the-art conventional, two branch and guided LLIE approaches. For the conventional LLIE, we compared the UFormer(UF) (Wang et al. 2022), EnlightenGAN(EnGAN) (Jiang et al. 2021), Restormer(RSTM) (Zamir et al. 2022), LLFormer(LLF) (Wang et al. 2023), and MIRNet (Zamir et al. 2020). For the two-branch LLIE, we compared the RUAS (Liu et al. 2021), DRBN (Yang et al. 2021a) and KinD (Zhang, Zhang, and Guo 2019). For the guided LLIE, we compared the RetinexNet(RN) (Wei et al. 2018) and SNR-Net (Xu et al. 2022).
Results Analysis
Quantitative Results
@@
To validate the quantitative performance of ECAFormer, we compared our model across seven public datasets using two metrics, PSNR and SSIM, with the results presented respectively in Table 1 and Table 2. The results on our collected Traffic-297 dataset are shown in Table 3. Specifically, our ECAFormer achieves a 3% improvement in PSNR and nearly a 1% improvement in SSIM over the suboptimal method. As we can observe, our ECAFormer demonstrates a 3% improvement in PSNR and nearly a 1% improvement in SSIM over the suboptimal method. The results indicate that our model achieves competitive performance while maintaining a relatively low parameter count and computational complexity.
UF RN EnGAN RUAS DRBN KinD RSTM LLF MIRNet SNR ECAFormer
LOL-v1 0.771 0.560 0.650 0.720 0.830 0.790 0.823 0.816 0.830 0.842 0.850
LOL-v2-r 0.771 0.567 0.617 0.723 0.831 0.641 0.827 0.792 0.82 0.849 0.853
LOL-v2-s 0.871 0.798 0.734 0.652 0.927 0.578 0.830 0.909 0.876 0.928 0.931
SID 0.577 0.578 0.543 0.581 0.577 0.583 0.649 - 0.605 0.625 0.660
SMID 0.792 0.684 0.674 0.744 0.781 0.634 0.758 - 0.762 0.805 0.810
SDSD-in 0.859 0.617 0.604 0.696 0.868 0.672 0.827 - 0.864 0.894 0.874
SDSD-out 0.748 0.629 0.616 0.743 0.841 0.654 0.802 - 0.837 0.866 0.862
Average 0.769 0.633 0.634 0.694 0.807 0.650 0.788 - 0.799 0.829 0.834
@@
Table 2: Comparison of SSIM across 7 public datasets using different methods
RN ZeroDCE++ EnGAN SNR ECAFormer
PSNR 18.56 18.35 18.19 30.06 33.26
SSIM 0.770 0.884 0.623 0.965 0.978
Table 3: Comparison of PSNR and SSIM on the traffic-297 dataset
Quality Results
@@
To validate the quality performance of ECAFormer, we present the visual comparison of the results, as shown in Figure 5-7. As we can observe, the recent fully supervised learning method, SNRNet, did not retain sufficient detail and exhibited blurring. ZeroDCE++, which adjusts the pixel curves, introduced excessive noise. In the detailed comparison illustrated in Figure 6, our method effectively preserves texture details while considering global brightness balance, indicating its ability to retain details and suppress noise. In Figure 8, we showcase the performance of our model across different scenes, including color discrepancies caused by nighttime shooting. Specifically, our method was tested in complex scenarios from SDSD-in, SDSD-out, and SMID datasets, with results indicating strong performance of ECAFormer. In Figure 7, we evaluated the performance in traffic scenes, specifically testing models trained on synthesized datasets under real-world nighttime scenarios. The complexity of lighting conditions in nighttime traffic scenes poses a significant challenge to models. While other methods may introduce noise and lose details, our model demonstrates exceptional visual performance by effectively preserving details while considering global illumination. In comparison, ZeroDCE++ exhibits overexposure issues, highlighting the superior visual results of our proposed model. These results indicate that our model adapts well to variations and produces high-quality outcomes closely approximating the ground truth, demonstrating the effectiveness of ECAFormer as proposed in this paper.
@@
Ablation Study
To assess the efficacy of the different modules, including 1) Visual-Semantic Feature (VSF), 2) Loss, and 3) DMSA, we made adjustments to the model structure and presented the PSNR results in given Table 4 and Figures 9 - 10. The results presented in Table 4 suggest that all three components positively impact the model’s performance and are effective. Among them, the DMSA module shows the most significant improvement in feature fusion.
@@
Table 4: Ablation study results on SDSD-out
Furthermore, in ECAFormer, we utilize the DMSA module to effectively manage cross-layer residual information, thereby minimizing the loss of details resulting from downsampling. This validates the capability of CSDMSA module to integrate residual information, as depicted in Figure 9. It is evident that without the CSDMSA module, the enhanced images suffer from blurring and significant loss of image detail. This outcome indicates that the CSDMSA module effectively mitigates the loss of details caused by downsampling by utilizing cross-layer residual information.
@@ Conclusion
In this paper, we propose a transformer-based method for LLIE, specifically utilizing DMSA blocks to enhance low-light images. Furthermore, the ECAFormer network combines the advantages of local detailed feature information extraction in CNNs and global information extraction in transformer networks. Therefore, ECAFormer is capable of extracting both global and local information from the image. Extensive experimental results validate the competitive advantages of ECAFormer compared to other state-of-the-art algorithms. In our future research, we will focus on refining the training approach to enable the model to effectively utilize disparate datasets and integrate additional prior knowledge to enhance its feature extraction capabilities. This is particularly important as ECAFormer currently relies on paired images, which significantly complicates the process of dataset collection.
Checklist
  • This paper:
    • Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes/partial/no/NA)
      Yes.
    • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes/no)
      Yes.
    • Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper (yes/no)
      Yes.
  • Does this paper make theoretical contributions? (yes/no)
    Yes.
    If yes, please complete the list below.
    • All assumptions and restrictions are stated clearly and formally. (yes/partial/no)
      Yes.
    • All novel claims are stated formally (e.g., in theorem statements). (yes/partial/no)
      Yes.
    • Proofs of all novel claims are included.(yes/partial/no)
      Partial.
    • Proof sketches or intuitions are given for complex and/or novel results. (yes/partial/no)
      Yes.
    • Appropriate citations to theoretical tools used are given. (yes/partial/no)
    • All theoretical claims are demonstrated empirically to hold. (yes/partial/no/NA)
      Yes.
    • All experimental code used to eliminate or disprove claims is included. (yes/no/NA)
      Yes.
  • Does this paper rely on one or more datasets? (yes/no)
    Yes
    If yes, please complete the list below.
    • A motivation is given for why the experiments are conducted on the selected datasets (yes/partial/no/NA)
      Yes.
    • All novel datasets introduced in this paper are included in a data appendix. (yes/partial/no/NA)
      Yes.
    • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no/NA)
      Yes.
    • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes/no/NA)
      Yes.
    • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes/partial/no/NA)
      Yes.
    • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (yes/partial/no/NA)
      Yes
  • Does this paper include computational experiments? (yes/no)
    Yes
    If yes, please complete the list below
    • Any code required for pre-processing data is included in the appendix. (yes/partial/no).
      Yes.
    • All source code required for conducting and analyzing the experiments is included in a code appendix. (yes/partial/no)
      Yes.
    • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no)
      Yes.
    • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes/partial/no)
      Partial.
    • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (yes/partial/no/NA)
      Yes.
    • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (yes/partial/no)
      Yes.
    • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes/partial/no)
      Yes.
    • This paper states the number of algorithm runs used to compute each reported result. (yes/no)
      Yes.
    • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes/no)
      Yes.
    • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes/partial/no)
      Yes.
    • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes/partial/no/NA)
      Partial. They can be found in the config file of the code that will be published later.
    • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes/partial/no/NA)
      Partial.

### v3

@@ 4 Experiment
4.1 Experimental Configurations
4.1.1 Datasets
@@
We conducted comparisons across multiple public datasets: LOL-v1 [19], LOL-v2 [30], SID [31] and SMID [32]. Additionally, we have collected the Traffic-297 dataset, which includes various traffic scenes. It contains an amount of real nighttime traffic scenes with complex lighting environments, which pose a significant challenge to low-light enhancement models.
4.1.2 Explementation Details
@@
We implemented the ECAFormer model using PyTorch. The model was trained on an NVIDIA RTX 4090 with 24 GB of VRAM using the Adam optimizer ($\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999), across a total of 250,000 iterations. The initial learning rate was set at 2 $\times$ 10-4 and gradually reduced to 1 $\times$ 10-6 through a cosine annealing schedule during the training process. Training samples were generated by randomly cropping 256 $\times$ 256 patches from pairs of low-/normal-light images. The batch size was set to 8. The training data was augmented with random rotations and flips to enhance variability and robustness. The training objective was to minimize both $\mathcal{L}_{p}$ 3.3.1 and $\mathcal{L}_{c}$ 3.3.2 between the enhanced images and their corresponding ground truths, ensuring high fidelity in the restoration of image details and color accuracy. To evaluate the performance, we employ the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) [33] as the primary metrics for evaluation.
4.1.3 Comparison Methods
@@
To validate the effectiveness of our derived ECAFormer, we conducted extensive comparisons with several state-of-the-art conventional, two branch and guided LLIE approaches. For the conventional LLIE, we compared the UFormer(UF) [7], EnlightenGAN(EnGAN) [21], Restormer(RSTM) [34], LLFormer(LLF) [35], and MIRNet [36]. For the two-branch LLIE, we compared the RetinexNet(RN) [19] and KinD [37]. For the guided LLIE, we compared the RUAS [20], DRBN [38], RetinexFormer(RF)[28], U2E-Net[39] and SNR-Net [27]
4.2 Results Analysis
4.2.1 Quantitative Results
@@
To validate the quantitative performance of ECAFormer, we compared our model across five public datasets using two metrics, PSNR and SSIM, with the results presented respectively in Table 1 and Table 2. The results on our collected Traffic-297 dataset are shown in Table 3. Specifically, our ECAFormer achieves a 3.7% improvement in PSNR and nearly a 1.3% improvement in SSIM over the suboptimal method. The results indicate that our model achieves competitive performance while maintaining a relatively low parameter count and computational complexity.
@@
Table 2: Comparison of SSIM of different methods across 5 public datasets.
Methods LOL-v1 LOLv2-r LOLv2-s SID SMID AVG Params
RN(BMVC 2018) 0.560 0.567 0.798 0.578 0.684 0.6374 0.84
KinD(ACMMM 2019) 0.790 0.641 0.578 0.583 0.634 0.6452 8.02
EnGAN(TIP 2021) 0.650 0.617 0.734 0.543 0.674 0.6436 114.35
RUAS(CVPR 2021) 0.720 0.723 0.652 0.581 0.744 0.6840 0.003
DRBN(TIP 2021) 0.830 0.831 0.927 0.577 0.781 0.7892 5.27
UF(CVPR 2022) 0.771 0.771 0.871 0.577 0.792 0.7564 5.29
RSTM(CVPR 2022) 0.823 0.827 0.830 0.649 0.758 0.7774 26.13
MIRNet(TPAMI 2022) 0.830 0.820 0.876 0.605 0.762 0.7786 31.76
SNR(CVPR 2022) 0.842 0.849 0.928 0.625 0.805 0.8098 4.01
LLF(AAAI 2023) 0.816 0.792 0.909 \ \ \ 24.55
RF(ICCV 2023) 0.816 0.814 0.930 0.673 0.815 0.8096 1.61
U2E-Net(PR 2024) 0.847 \ \ \ \ \ \
ECAFormer 0.850 0.853 0.931 0.660 0.810 0.8208 2.50
Table 3: Comparison of PSNR and SSIM on the traffic-297 dataset.
RN ZeroDCE++ EnGAN SNR RF ECAFormer
PSNR 18.56 18.35 18.19 30.06 32.17 33.26
SSIM 0.770 0.884 0.623 0.965 0.975 0.978
4.2.2 Quality Results
@@
To validate the quality performance of ECAFormer, we present the visual comparison of the results, as shown in Figure 5-8. As we can observe, the recent fully supervised learning method, SNRNet, did not retain sufficient detail and exhibited blurring. ZeroDCE++, which adjusts the pixel curves, introduced excessive noise. In the detailed comparison illustrated in Figure 6, our method effectively preserves texture details while considering global brightness balance, indicating its ability to retain details and suppress noise. In Figure 7, we showcase the performance of our model across different scenes, including color discrepancies caused by nighttime shooting. Specifically, our method was tested in complex scenarios from SDSD-in [40], SDSD-out, and SMID datasets, with results indicating strong performance of ECAFormer. In Figure 8, we evaluated the performance in traffic scenes, specifically testing models trained on synthesized datasets under real-world nighttime scenarios. The complexity of lighting conditions in nighttime traffic scenes poses a significant challenge to models. While other methods may introduce noise and lose details, our model demonstrates exceptional visual performance by effectively preserving details while considering global illumination. In comparison, ZeroDCE++ exhibits overexposure issues, highlighting the superior visual results of our proposed model. These results indicate that our model adapts well to variations and produces high-quality outcomes closely approximating the ground truth, demonstrating the effectiveness of ECAFormer as proposed in this paper.
@@
4.3 Ablation Study
To assess the efficacy of the different modules, including (1) Visual-Semantic Feature (VSF): remove the visual semantic module and directly pass the input to DMSA., (2) Loss: only use $L_{1}$ loss without perceptual loss and charbonnier loss, and (3) DMSA: replace DMSA with conventional MHSA, we made adjustments to the model structure and presented the PSNR results in given Table 4 and Figures 9 - 10. The results presented in Table 4 suggest that all three components positively impact the model’s performance and are effective. Among them, the DMSA module shows the most significant improvement in feature fusion.
Table 4: Ablation study results on SDSD-out.
@@ 4.3 Ablation Study
Furthermore, in ECAFormer, we utilize the DMSA module to effectively manage cross-layer residual information, thereby minimizing the loss of details resulting from downsampling. This validates the capability of the CSDMSA module to integrate residual information, as depicted in Figure 9. It is evident that without the CSDMSA module, the enhanced images suffer from blurring and significant loss of image detail. This outcome indicates that the CSDMSA module effectively mitigates the loss of details caused by downsampling by utilizing cross-layer residual information.
@@ 5 Conclusion
In this paper, we propose a transformer-based method for LLIE. Specifically, we use DMSA blocks to enhance low-light images by enhancing the semantic and visual components of the image. Furthermore, the ECAFormer network combines the advantages of local detailed feature information extraction in CNNs and global information extraction in transformer networks, which effectively reveals the local and global information of the image. Thus, the detailed information can be preserved after enhancement. Extensive experimental results validate the competitive advantages of ECAFormer compared to other state-of-the-art algorithms. In our future research, we will focus on refining the training approach to enable the model to effectively utilize disparate datasets and integrate additional prior knowledge to enhance its feature extraction capabilities.
ACKNOWLEDGMENTS
This work was partially supported by the National Natural Science Foundation of China (62306051,6230073196,62481540175,), the Fundamental Research Funds for the Central Universities (NJ2022028), and the Scientific and the Technological Research Program of Chongqing Municipal Education Commission (KJQN202300718). Weikai Li was partly supported by the National Natural Science Foundation of China (62076124).